OpenStudy (anonymous):

last question :) A 60 foot light pole casts a shadow that is 40 feet long. At the same time, a nearby light post casts a shadow that is 15 feet long. How tall is the light post?

jimthompson5910 (jim_thompson5910):

(Height of light pole)/(Length of light pole's shadow) = (Height of light post)/(Length of light post's shadow) 60/40 = x/15 60*15 = 40x 900 = 40x 900/40 = x 22.5 = x x = 22.5 So the height of the light post is 22.5 feet.

jimthompson5910 (jim_thompson5910):

You start by setting up a proportion. Then you cross multiply. From there, you divide both sides by 40 to solve for x.

OpenStudy (anonymous):

thank you sooooo very much :)

jimthompson5910 (jim_thompson5910):

you're welcome